Sadik Industries must install $1 million of new machinery in its Texas plant. It can obtain a bank loan for 100% of the required amount. Alternatively, a Texas investment banking firm that represents a group of investors believes it can arrange for a lease financing plan. Assume that the following facts apply.

(1) The equipment falls in the MACRS 3 -year class.
(2) Estimated maintenance expenses are $50,000 per year.
(3) The firm's tax rate is 34%.
(4) If the money is borrowed, the bank loan will be at a rate of 14%, amortized in three equal installments at the end of each year.
(5) The tentative lease terms call for payments of $320,000 at the end of each year for 3 years. The lease is a guideline lease.
(6) Under the proposed lease terms, the lessee must pay for insurance, property taxes, and maintenance.
(7) Sadik must use the equipment if it is to continue in business, so it will almost certainly want to acquire the property at the end of the lease. If it does, then under the lease terms it can purchase the machinery at its fair market value at Year 3. The best estimate of this market value is $200,000, but it could be much higher or lower under certain circumstances.

If purchased at Year 3, the used equipment would fall into the MACRS 3-year class. Sadik would actually be able to make the purchase on the last day of the year (i.e., slightly before Year 3), so Sadik would get to take the first depreciation expense at Year 3 (the remaining depreciation expenses would be at Year 4 through Year 6). On the time line, Sadik would show the cost of the used equipment at Year 3 and its depreciation expenses starting at Year 3.

To assist management in making the proper lease-versus-buy decision, you are asked to answer the following.

a. What is the net advantage of leasing? Should Sadik take the lease?

b. Consider the $200,000 estimated residual value. How high could the residual value get

before the net advantage of leasing falls to zero?

c. The decision almost can be considered a bet on the future residual value. Do you think

the residual cash flows are equal in risk to the other cash flows? If not, how might you address

this issue? ( Hint:   if you discount a negative cash flow at a higher rate, you get a better NPV

� the NPV of a negative cash flow stream is less negative at high discount rates.)
Answer #1
PURCHASE OPTION
Equipment Cost $1,000,000
3 year MACRS
Equipment Cost=$1000000 $1,000,000 TAX RATE=34%
A B=A*$1000000 C D=B*34%
Depreciation Amount of Accumulated Depreciation
Year Rate Depreciation Depreciation Tax Shield
1 33.33% $333,300 $333,300 $69,993
2 44.45% $444,500 $777,800 $93,345
3 14.81% $148,100 $925,900 $31,101
4 7.41% $74,100 $1,000,000 $15,561
After tax cost of debt =14*(1-0.34) 9.24%
Discount Rate =9.24%=0.0924
INTEREST AND PRINCIPAL REPAYMENT ON AMOUNT BORROWED
Pv Amount Borrowed = $1,000,000
Nper Number of years of repayment                              3
Rate Interest Rate 14%
PMT Annual repayment in three equal instalments $430,731 (Using PMT function of excelwith Rate=14%,Nper=3, Pv=-1000000)
REPAYMENT SCHEDULE
Year 1 2 3
A Beginning Balance $1,000,000 $709,269 $377,835
B Amount of annual payment $430,731 $430,731 $430,731
C=A*14% Interest $140,000 $99,298 $52,897
D=B-C Principal $290,731 $331,434 $377,835
E=A-D Ending Balance $709,269 $377,835 $0
F=C*34% Interest Tax Shield $47,600 $33,761 $17,985
N Year 0 1 2 3 4 5 6
Annual Cash Inflows:
.(1) Cash Flow for annual repayment -$430,731 -$430,731 -$430,731 $0 $0 $0
.(2) Depreciation Tax shield $69,993 $93,345 $31,101 $15,561 $0 $0
.(3) Interest Tax Shield $47,600 $33,761 $17,985 $0 $0 $0
CF=(1)+(2)+(3) Net Cash Flow -$313,138 -$303,625 -$381,646 $15,561 $0 $0 SUM
PV=CF/(1.0924^N) Present Value of Cash Flow at discount rate9.24% -$286,652 -$254,434 -$292,762 $10,927 $0 $0 -$822,921
NET PRESENT VALUE(NPV) -$822,921
LEASE OPTION
After tax lease payment =320000*(1-0.34) $211,200
N Year 0 1 2 3 4 5 6
.(1) After Tax Lease Payment -$211,200 -$211,200 -$211,200 $0 $0 $0
.(2) Terminal Payment ($200,000)
a MACRS Depreciation rate 33.33% 44.45% 14.81% 7.41%
b=a*200000 Annual Depreciation $66,660 $88,900 $29,620 $14,820
c=b*34% Depreciation Tax shield $22,664 $30,226 $10,071 $5,039
CF=(1).+(2)+c Net Cash Flow -$211,200 -$211,200 -$388,536 $30,226 $10,071 $5,039 SUM
PV=CF/(1.0924^N) Present Value of Cash Flow at discount rate9.24% -$193,336 -$176,983 -$298,048 $21,225 $6,474 $2,965 -$637,702
NET PRESENT VALUE(NPV) -$637,702
LEASE OPTION IS RECOMMENDED
NPV of Cost is Lower
Net Advantage of Leasing=822921-637702 $185,219

H18 - X for =PMT(H17,H16,-H15) $1,000,000 PURCHASE OPTION Equipment Cost 3 year MACRS Equipment Cost=$1000000 $1,000,000 Depr4 2 $211,200 6 $0 -$211,200 $0 $0 LEASE OPTION After tax lease payment =320000*(1-0.34) $211,200 Year After Tax Lease Payment